OpenStudy (anonymous):

2/3(y+2) - 5/6(y-1) = 5/4(y-3) - 4 solve for y?

OpenStudy (anonymous):

This is how i do these. find a lcm and multiply all your fraction by it inorder to clear the fractions. then you will distribute into the quanities. transpose varibles and constants and divde by the coeficient infront of your variable.

OpenStudy (anonymous):

So after clearing the fractions you end up with 8(y+2)-10(y-1)=15(y-3)

OpenStudy (anonymous):

8y+16-10y+10=15y-45

OpenStudy (anonymous):

simplify boths sides 16-2y=15y-45

OpenStudy (anonymous):

transpose 61=17y

OpenStudy (anonymous):

divide both sides by 17 you end up with 61/17=y

OpenStudy (anonymous):

I think I understand it now...Thanks for the little trick idea! Ill try that in future problems!

OpenStudy (anonymous):

Just remember if you do it to a fraction to clear it you must do it to all the remaining quanties in the equation unless they are in perenthises